can someone help me pls?
solve.
if y varies directly as x, and x=12 when y=5, find x when y=3

Answer by ewatrrr(24785)

 
Hi,
y varies directly as x OR
y = kx |y=5 when x=12
5 = k*12
5/12 = k
y =(5/12)x | y = 3 THEN x = (12/5)*3 = 36/5